the carpenter saws a board that is 22 feet long into two pieces. One piece is to be 1 foot longer than twice the length of the shorter piece. Find the length of each piece.
Answer by fractalier(6550)

Call the length of the shorter piece x. Then the longer piece will be 1 + 2x.
Together they must be 22 feet...the algebra looks like this:
x + (1 + 2x) = 22
3x + 1 = 22
3x = 21
x = 7
2x + 1 = 15
Your pieces are 7 and 15 feet long.